# ControlMM: Controllable Masked Motion Generation

### v1

@@ 5.1 Quantitative Comparison to State-of-the-art Approaches
colspec = lcccccccc, cell42 = r=6, cell112 = r=3, cell142 = r=3, cell172 = r=3, cell202 = r=3, cell232 = r=3, cell262 = r=3, cell292 = r=2, cell312 = r=2, hline1-3, 10, 14, 17, 20, 23, 26, 29, 31, 33 = -, Method & Joint R-Precision Top-3 $\uparrow$ FID $\downarrow$ Diversity $\rightarrow$ Foot Skating Ratio $\downarrow$ Traj. Err. (50 cm) $\downarrow$ Loc. Err. (50 cm) $\downarrow$ Avg. Err. $\downarrow$
@@
ControlMM 0.811 0.049 9.533 0.0545 0.0000 0.0000 0.0126
GMD (Karunratanakul et al., 2023) only addresses the pelvis location on the ground plane (xz coordinates). To ensure a fair comparison, we follow OmniControl (Xie et al., 2023) and compare GMD in managing the full 3D position of the pelvis (xyz coordinates). The first section of Table 1 resents results for models trained on the pelvis alone to ensure a fair comparison with previous state-of-the-art methods on the HumanML3D (Guo et al., 2022b) dataset. $\rightarrow$ mean means closer to real data is better. Our model demonstrates significant improvements across all evaluation metrics. When compare to TLControl, the FID score notably decreased from 0.271 to 0.061, indicating superior generation quality. Additionally, the R-Precision increased from 0.779 to 0.809, highlighting improved alignment between the text descriptions and the generated motions. In terms of spatial control accuracy, both Trajectory Error and Location Error dropped to zero, while the average error decreased to 0.91 cm, indicating highly precise spatial control. Furthermore, our model outperforms existing methods in both Diversity and Foot Skating Ratio metrics. In the second section, Train on All Joints, we follow the evaluation from OmniControl (Xie et al., 2023), as our model supports control of any joint, not just the root (pelvis). We train the model to control multiple joints, specifically the pelvis, left foot, right foot, head, left wrist, and right wrist. The Cross experiment shows 63 cross-joint combinations (details in Appendix. A.11), while Average reflects the average performance across each joint. Our model outperforms all other methods across all joint configurations, including Average and Cross. Compared to OmniControl, our model delivers superior quality in Cross, evidenced by a FID score drop to 0.049 and an R-Precision increase to 0.811. In contrast, OmniControl struggles with multiple joints, as its FID score spikes to 0.624—almost triple its performance on the pelvis alone. Moreover, our model maintains zero Trajectory and Location Errors, while preserving Diversity, whereas OmniControl’s Trajectory Error increase to 0.2147 and Diversity significantly drops to 9.016, indicating our model’s robust handling of multiple control signals.
5.2 Qualitative Comparison to State-of-the-art Approaches
We visualize the generated motion using GMD (Karunratanakul et al., 2023) and OmniControl (Xie et al., 2023) in Fig. 5. The motion is generated based on the prompt “a person walks forward and waves his hands,” with the pelvis and right wrist controlled in a zigzag pattern. Since GMD can only control the pelvis, we apply control only to the pelvis for GMD. However, it fails to follow the zigzag pattern, tending instead to move in a straight line. OmniControl receives control signals for both the pelvis and right wrist. Yet, it not only fails to follow the root trajectory (pelvis) but also does not adhere to the zigzag pattern for the right wrist. In contrast, our ControlMM demonstrates realistic motion with precise spatial control for both the pelvis and the right wrist, accurately following the intended zigzag pattern.
Figure 5: Visualization comparisons to state-of-the-art methods. The plots on the top display the top view of pelvis control (root trajectory), while the bottom plot shows the side view of the right wrist. Red represents the control signal, and Blue represents the generated joint motion.
5.3 Body Part Editing
With spatial signal control, our model is capable of conditioning on multiple joints, which can be treated as distinct body parts, while generating the remaining body parts based on text input. In Table 2 We quantitatively compare our approach to existing methods designed for this task, including MDM (Tevet et al., 2022b) and MMM (Pinyoanuntapong et al., 2024b). Additionally, we compare it with OmniControl (Xie et al., 2023), which also supports spatial signal control. However, our evaluation demonstrates that OmniControl performs poorly in this task. Following the evaluation protocol from (Pinyoanuntapong et al., 2024b), we condition the lower body parts on ground truth for all frames and generate the upper body based on text descriptions using the HumanML3D dataset (Guo et al., 2022b). Our model is evaluated without retraining, using the same model as in the Train on All Joints setup, ensuring a fair comparison with OmniControl, which is trained on a subset of joints. Specifically, we condition only on the pelvis, left foot, and right foot as the lower body signals.
The results show that MDM struggles significantly when conditioned on multiple joints, with the FID score increasing to 4.827. Although OmniControl supports multiple joint control, our experiments reveal that it also suffers under these conditions, with its FID score rising to 1.213. This is consistent with the Cross-Joint evaluation in Table 1, which evaluate on multiple joint combination, where OmniControl’s FID score deteriorates considerably. MMM performs well in this task but requires retraining with separate codebooks for upper and lower body parts. In contrast, our model outperforms all other methods across all metrics without any retraining. When comparing to the ‘Train on Pelvis Only’ setup in Table 1, our model achieves similar FID and R-Precision scores, highlighting its robustness in handling multiple joint control signals.
Top1 Top2 Top3
MDM (Tevet et al., 2022b) 0.298 0.462 0.571 4.827 4.598 7.010
OmniControl (Xie et al., 2023) 0.374 0.550 0.656 1.213 5.228 9.258
MMM (Pinyoanuntapong et al., 2024b) 0.500 0.694 0.798 0.103 2.972 9.254
ControlMM (ours) 0.517 0.708 0.804 0.074 2.945 9.380
6 Ablation Study
6.1 Qualitative results
We visualize each component in Fig. 6 by controlling the pelvis and left wrist with the text prompt “a person walks in a circle counter-clockwise.” (a) Motion Control Model: The overall motion is realistic but the controlled joints (pelvis and left wrist) deviate significantly from the spatial control signals. (b) Logits Editing: The root positions (pelvis) are closer to the spatial control signal, but the left wrist positions remain inaccurate. (c) Embedding Editing: Both the pelvis and left wrist positions align more closely with the spatial control signals, but the motion lacks realism because Embedding Editing only adjusts the motion at the end of the generation process. (d) Full Model: With all components active, the model generates realistic motion with high precision to match the control signals, while OmniControl fails to follow the control signals for both the pelvis and left wrist.
Figure 6: Qualitative comparisons of each component and the baseline
6.2 Component Analysis
Table 3: Ablation results of all combinations of the main components.
{tblr}
colspec = lccccccccc, hline1,9,10 = -, # & Logits Editing. Emb Editing. Motion Control Model R-Precision Top-3 $\uparrow$ FID $\downarrow$ Diversity $\rightarrow$ Foot Skating Ratio $\downarrow$ Traj. Err. (50 cm) $\downarrow$ Loc. Err. (50 cm) $\downarrow$ Avg. Err. $\downarrow$
1 ✗ ✗ ✗ 0.807 0.095 9.672 0.0527 0.5066 0.3511 0.6318
2 ✗ ✗ 0.813 0.105 9.615 0.0529 0.2323 0.1175 0.2361
3 ✗ 0.786 0.190 9.294 0.0616 0.0063 0.0005 0.0283
4 ✗ 0.795 0.142 9.402 0.0577 0.0032 0.0002 0.0218
5 ✗ ✗ 0.802 0.128 9.475 0.0594 0.3914 0.2400 0.4041
6 0.814 0.051 9.557 0.0541 0.1302 0.0623 0.1660
7 0.806 0.069 9.425 0.0568 0.0005 0.0000 0.0124
8 0.809 0.061 9.496 0.0547 0.0000 0.0000 0.0098
The key components of our model are Logits Editing, Codebook Editing, and Motion Control Model. To understand how each component impact the quality and spatial control error. We conduct ablation experiments using same evaluation as Table 1.
From Table 6.2, Without any control (#1), the model achieves the highest diversity and the lowest Foot Skating Ratio, indicating strong realism in the generated motion. The FID score is also on par. However, all spatial errors are poor due to the absence of spatial control components in the model. For the experiments without Embedding Editing (#1, #2, #5, #6), both FID scores and R-Precision are notable, particularly in #6, which combines Logits Editing and the Motion Control Model to enhance generation quality. In contrast, #3, which solely uses Embedding Editing, exhibits the worst FID score and Foot Skating Ratio while showing acceptable spatial control errors. This experiment highlights that while Embedding Editing can reduce generation errors, it may negatively impact the overall quality. Conversely, incorporating Logit Editing and Motion Control Model during each iteration improves both quality and spatial control errors, as demonstrated in #8.
6.3 Density of Spatial Control Signal
Table 4: Ablation results on different densities.
{tblr}
colspec = lccccccc, hline1,7,8 = -, Density R-Precision Top-3 $\uparrow$ FID $\downarrow$ Diversity $\rightarrow$ Foot Skating Ratio $\downarrow$ Traj. Err. (50 cm) $\downarrow$ Loc. Err. (50 cm) $\downarrow$ Avg. Err. $\downarrow$
1 0.804 0.077 9.526 0.0551 0.0000 0.0000 0.0010
2 0.806 0.087 9.475 0.0553 0.0000 0.0000 0.0034
5 0.811 0.078 9.499 0.0553 0.0000 0.0000 0.0098
49 (50%) 0.812 0.055 9.507 0.0536 0.0001 0.0000 0.0168
196 (100%) 0.814 0.054 9.514 0.0543 0.0002 0.0000 0.0164
In table 6.3, we provide a detailed analysis of ControlMM’s performance across five different spatial control density levels, where the model is trained for pelvis control using the HumanML3D dataset. The results show that increasing the spatial control improves the quality: the FID score decreases from 0.077 with 1-frame control to 0.054 with full 196-frame (100%) control. Similarly, R-Precision improves from 0.804 at 1-frame density to 0.814 at 196-frame (100%) density. However, the Average Error shows the opposite trend—more spatial control leads to higher error, as the model is required to target more specific points.
7 Conclusion
In this work, we present ControlMM, a new method that incorporates spatial control signals into the Masked Motion Model. ControlMM is the first model that enables precise control over quantized motion tokens while maintaining high-quality motion generation at faster speeds, consistently outperforming diffusion-based controllable frameworks. ControlMM introduces two key innovations: Masked Consistency Modeling uses random masking and reconstruction to ensure that the generated motions are of high fidelity, while also reducing inconsistencies between the input control signals and the motions produced. Inference-Time Logit and Codebook Editing fine-tunes the predicted motion distribution to better match the input control signals, enhancing precision and making ControlMM adaptable for various tasks. ControlMM has a wide range of applications, including any joint any frame control, obstacle avoidance, and body part timeline control.
Appendix A Appendix
A.1 Overview
  • Section A.2: Implementation Details
  • Section A.3: Inference speed, quality, and errors Details
  • Section A.4: Speed of each component
  • Section A.5: Quantitative result for all joints of ControlMM-Fast
  • Section A.6: Ablation on less number of generation iteration
  • Section A.7: Analysis of Logits Editing and Motion Control Model
  • Section A.8: Approximated Mask Embedding for Decoder
  • Section A.9: Body Part Timeline Control
  • Section A.10: Dual-Space Categorical Straight-Through Estimator
  • Section A.11: Cross Combination
Video visualization can be found at https://exitudio.github.io/ControlMM-page/
A.2 Implementation Details
We adopt the pretrained MoMask (Guo et al., 2023) model as a base model and maintain the default hyperparameter settings. To improve robustness to text variation, we randomly drop 10% of the text conditioning, which also allows the model to be used for Classifier-Free Guidance (CFG). The weight for Eq. 5 is set to $\alpha=0.1$. We use a codebook of size 512, with embeddings of size 512 and 6 residual layers. The Transformer embedding size is set to 384, with 6 attention heads, each with an embedding size of 64, distributed across 8 layers. This configuration demonstrates the feasibility of converting between two different embedding sizes and spaces using the Dual-Space Categorical Straight-Through Estimator. The encoder and decoder downsample the motion sequence length by a factor of 4 when mapping to token space. The learning rate follows a linear warm-up schedule, reaching 2e-4 after 2000 iterations, using AdamW optimization. The mini-batch size is set to 512 for training RVQ-VAE and 64 for training the Transformers. During inference, the CFG scale is set to $cfg=4$ for the base layer and $cfg=5$ for the 6 layers of residual, with 10 iterations for generation. We use pretrained CLIP model (Radford et al., 2021) to generate text embeddings, which have a size of 512. These embeddings are then projected down to a size of 384 to match the token size used by the Transformer. Motion Control Model is a trainable copy of Masked Transformer with the zero linear layer connect to the output each layer of the Masked Transformer. During inference, Logits and Codebook Editing applies L2 loss with a learning rate of $0.06$ for 100 steps in Codebook Editing for each of the 10 generation iterations and 600 steps in Logits Editing. We apply temperature of 1 for all 10 iterations and 1e-8 for residual layers.
A.3 Inference speed, quality, and errors
@@
We compare the speed of three different configurations of our model against state-of-the-art methods as shown in Table A.3. The first setting, ControlMM-Fast, uses 100 steps of Codebook Editing without Logits Editing. This setup achieves results comparable to OmniControl, but is over 20 times faster. It also slightly improves the Trajectory and Location Errors, while the FID score is only 25% of OmniControl’s, indicating high generation quality. The second setting, ControlMM-Medium, increases the Codebook Editing to 600 steps, which further improves accuracy. The Location Error is reduced to zero, although the FID score slightly worsens. Lastly, the ControlMM-Accurate model, which is the default setting used in other tables in this paper, uses 600 steps of Codebook Editing and 100 steps of Logits Editing. This configuration achieves extremely high accuracy, with both the Trajectory and Location Errors reduced to zero and the Average Error below 1 cm (0.0098 meters). Importantly, these settings can be adjusted during inference without retraining the model, making them suitable for both real-time and high-performance applications.
@@
colspec = lccccccc, hline1,6 = -, Model Speed $\downarrow$ R-Precision Top-3 $\uparrow$ FID $\downarrow$ Diversity $\rightarrow$ Foot Skating Ratio $\downarrow$ Traj. Err. (50 cm) $\downarrow$ Loc. Err. (50 cm) $\downarrow$ Avg. Err. $\downarrow$
@@
ControlMM-Fast 4.94 s 0.808 0.0590 9.444 0.0570 0.0200 0.0075 0.0550
ControlMM-Medium 25.23 s 0.806 0.0690 9.425 0.0568 0.0005 0.0000 0.0124
ControlMM-Accurate 71.72 s 0.809 0.061 9.496 0.0547 0.0000 0.0000 0.0098
A.4 Speed of each component
We report the inference time for each component in Table 6, with all measurements taken on an NVIDIA A100. The Base model, which includes only the Masked Transformer with Residual layers and Decoder (without any spatial control signal module), has an inference time of 0.35 second. The Motion Control Model is highly efficient, requiring only 0.24 seconds for inference. The Codebook Editing and Logits Editing components take 24.65 seconds and 46.5 seconds, respectively. In total, the ControlMM-Accurate model has a generation time of 71.73 seconds. Note that this setting is using 100 steps of Codebook Editing for 10 iterations and 600 steps of Logits Editing.
@@ A.4 Speed of each component
Base Motion Control Model Codebook Editing Logits Editing Full
@@
A.5 Quantitative result for all joints of ControlMM-Fast
Table A.5 presents the evaluation results for ControlMM-Fast, which uses 100 steps of Codebook Editing without Logits Editing. This evaluation includes a “cross” assessment that evaluates combinations of different joints, as detailed in Section A.11. The results can be compared to those of the full model (ControlMM-Accurate) and state-of-the-art models shown in Table 1. Additionally, ”lower body” refers to the conditions involving the left foot, right foot, and pelvis, which allows for the evaluation of upper body editing tasks, as illustrated in Table 2.
Table 7: Quantitative result for all joints of ControlMM-Fast
@@ A.5 Quantitative result for all joints of ControlMM-Fast
A.6 Ablation on less number of generation iteration
In this section, we perform an ablation study on the number of iterations used in the generation process. Following the MoMask architecture (Guo et al., 2023), we adopt the same setting of 10 iterations for generation. However, the integration of Logits Editing and the Motion Control Model enhances the quality of the generated outputs with fewer iterations, as demonstrated in Table A.6. Notably, with just 1 iteration, the results are already comparable to those achieved by TLControl (Wan et al., 2023). Furthermore, after 4 iterations, the evaluation metrics are on par with those obtained after 10 iterations.
Table 8: Quantitative result for different number of iterations with Logits Editing and Motion Control Model
@@ A.6 Ablation on less number of generation iteration
colspec = lccccccc, hline1,8 = -, # of iterations R-Precision Top-3 $\uparrow$ FID $\downarrow$ Diversity $\rightarrow$ Foot Skating Ratio $\downarrow$ Traj. Err. (50 cm) $\downarrow$ Loc. Err. (50 cm) $\downarrow$ Avg. Err. $\downarrow$
@@
To further investigate the influence of Logits Editing and the Motion Control Model for lesser iterations, we remove these components and experiment with various numbers of iterations, as shown in Table A.6. Reducing the number of iterations significantly decreases the quality of the generated outputs, resulting in an FID score of 1.196 with only 1 iteration. Even with 10 iterations, the FID score remains at 0.190, highlighting the improvements by integrating Logits Editing and the Motion Control Model.
Table 9: Quantitative result for different number of iterations without Logits Editing and Motion Control Model
@@
colspec = lccccccc, hline1,8 = -, # of iterations R-Precision Top-3 $\uparrow$ FID $\downarrow$ Diversity $\rightarrow$ Foot Skating Ratio $\downarrow$ Traj. Err. (50 cm) $\downarrow$ Loc. Err. (50 cm) $\downarrow$ Avg. Err. $\downarrow$
@@
A.7 Analysis of Logits Editing and Motion Control Model
To understand the impact of Logits Editing and Motion Control Model on the generation process, we visualize the maximum probability for each token prediction from the Masked Transformer. The model predicts 49 tokens over 10 iterations. We show results both before and after applying Logits Editing, and with and without the Motion Control Model. The maximum probability can be expressed as the relative value of the logits corresponding to all codes in the codebook in the specific token position and iteration, as computed by the Softmax function. We visualize the output using the Softmax function instead of Gumbel-Softmax. By removing the Gumbel noise, Gumbel-Softmax reduces to a regular Softmax function:

### v2

@@ A.5 Full Evaluation on All Joint
colspec = lcccccccc, cell42 = r=7, cell112 = r=3, cell152 = r=3, cell182 = r=3, cell212 = r=3, cell242 = r=3, cell272 = r=3, cell302 = r=2, cell322 = r=2, hline1-3, 11, 15, 18, 21, 24, 27, 30, 32, 34 = -, Method & Joint R-Precision Top-3 $\uparrow$ FID $\downarrow$ Diversity $\rightarrow$ Foot Skating Ratio $\downarrow$ Traj. Err. (50 cm) $\downarrow$ Loc. Err. (50 cm) $\downarrow$ Avg. Err. $\downarrow$
@@
MaskControl 0.811 0.049 9.533 0.0545 0.0000 0.0000 0.0126
A.6 Inference speed, quality, and errors
@@
We compare the speed of three different configurations of our model against state-of-the-art methods as shown in Table A.6. The first setting, MaskControl-Fast, uses only 100 iterations of Logits Optimization in the last step of unmasking process. This setup achieves results comparable to OmniControl, but is over 20 times faster. It also slightly improves the Trajectory and Location Errors, while the FID score is only 25% of OmniControl’s, indicating high generation quality. The second setting, MaskControl-Medium, increases the Logits Optimization to 600 iterations, which further improves accuracy. The Location Error is reduced to zero, although the FID score slightly worsens. Lastly, the MaskControl-Accurate model, which is the default setting used in other tables in this paper, uses 600 iterations of Logits Optimization in the last step of unmasking process and 100 iterations of 1-9 step Logits Optimization in the last step of unmasking process. This configuration achieves extremely high accuracy, with both the Trajectory and Location Errors reduced to zero and the Average Error below 1 cm (0.98 cm). Importantly, these settings can be adjusted during inference without retraining the model, making them suitable for both real-time and high-performance applications. Fig. A.6 compares FID, Location Error, and speed.
@@
colspec = lccccccc, hline1,6 = -, Model & Speed $\downarrow$ R-Precision Top-3 $\uparrow$ FID $\downarrow$ Diversity $\rightarrow$ Foot Skating Ratio $\downarrow$ Traj. Err. (50 cm) $\downarrow$ Loc. Err. (50 cm) $\downarrow$ Avg. Err. $\downarrow$
@@
MaskControl-Fast 4.94 s 0.808 0.059 9.444 0.0570 0.0200 0.0075 0.0550
MaskControl-Medium 25.23 s 0.806 0.069 9.425 0.0568 0.0005 0.0000 0.0124
MaskControl-Accurate 71.72 s 0.809 0.061 9.496 0.0547 0.0000 0.0000 0.0098
A.7 Speed of each component
We report the inference time for each component in Table 9, with all measurements taken on an NVIDIA A100. The Base model, which includes only the Masked Transformer with Residual layers and Decoder (without any spatial control signal module), has an inference time of 0.35 second. The Logits Regularizer is highly efficient, requiring only 0.24 seconds for inference. The Logits Optimization takes 24.65 seconds for unmasking step 1-9 and 46.5 seconds in the last step, respectively. In total, the MaskControl-Accurate model has a generation time of 71.73 seconds. Note that this setting is using 100 iterations of Logits Optimization for 1-9 steps and 600 iterations of the last step.
@@ A.7 Speed of each component
Base Logits Regularizer Logits Optimization (1-9 steps) Logits Optimization (last step) Full
@@
A.8 Quantitative result for all joints of MaskControl-Fast
Table A.8 presents the evaluation results for MaskControl-Fast, which uses 100 iterations of Logits Optimization. This evaluation includes a “cross” assessment that evaluates combinations of different joints, as detailed in Section A.13. The results can be compared to those of the full model (MaskControl-Accurate) and state-of-the-art models shown in Table 2. Additionally, ”lower body” refers to the conditions involving the left foot, right foot, and pelvis, which allows for the evaluation of upper body editing tasks, as illustrated in Table 3.
Table 10: Quantitative result for all joints of MaskControl-Fast
@@ A.8 Quantitative result for all joints of MaskControl-Fast
A.9 Ablation on less number of generation step
In this section, we perform an ablation study on the number of steps used in the generation process. Following the MoMask architecture [16], we adopt the same setting of 10 steps for generation. However, the integration of Logits Optimization and the Logits Regularizer enhances the quality of the generated outputs with fewer steps, as demonstrated in Table A.9. Notably, with just 1 step, the results are already comparable to those achieved by TLControl [51]. Furthermore, after 4 steps, the evaluation metrics are on par with those obtained after 10 steps.
Table 11: Quantitative result for different number of steps
@@ A.9 Ablation on less number of generation step
colspec = lccccccc, hline1,8 = -, # of steps R-Precision Top-3 $\uparrow$ FID $\downarrow$ Diversity $\rightarrow$ Foot Skating Ratio $\downarrow$ Traj. Err. (50 cm) $\downarrow$ Loc. Err. (50 cm) $\downarrow$ Avg. Err. $\downarrow$
@@
To further investigate the influence of Logits Optimization and the Logits Regularizer for lesser steps, we remove these components and experiment with various numbers of steps and apply Logits Regularizer only last step, as shown in Table A.9. Reducing the number of steps significantly decreases the quality of the generated outputs, resulting in an FID score of 1.196 with only 1 step. Even with 10 steps, the FID score remains at 0.190, highlighting the improvements by integrating Logits Optimization and the Logits Regularizer.
Table 12: Quantitative result for different number of steps without LogitsOptimization and Logits Regularizer
@@
colspec = lccccccc, hline1,8 = -, # of steps R-Precision Top-3 $\uparrow$ FID $\downarrow$ Diversity $\rightarrow$ Foot Skating Ratio $\downarrow$ Traj. Err. (50 cm) $\downarrow$ Loc. Err. (50 cm) $\downarrow$ Avg. Err. $\downarrow$
@@
A.10 Analysis of Logits Optimization and Logits Regularizer
To understand the impact of Logits Optimization and Logits Regularizer on the generation process, we visualize the maximum probability for each token prediction from the Masked Transformer. The model predicts 49 tokens over 10 steps. We show results both before and after applying Logits Optimization, and with and without the Logits Regularizer. The maximum probability can be expressed as the relative value of the logits corresponding to all codes in the codebook in the specific token position and step, as computed by the Softmax function. We visualize the output using the Softmax function instead of Gumbel-Softmax. By removing the Gumbel noise, Gumbel-Softmax reduces to a regular Softmax function: